Reports 1-1 of 1 Clear search Modify search
On-call intervention (General)
derossi, spinicelli, chiummo - 13:12 Monday 17 May 2021 (51810) Print this report
Comment to On-call intervention (51796)

This morning we were called by Berni due to frequent IMC unlocks.

As an attempt to mitigate the problem we performed the following actions:

- we moved the electronics of the ML away from the same breadbord of the optical setup (1st attached picture).

- we tuned the input ML polarization by turning the waveplate before the collimator from 182° to 185° (2nd plot) while checking the stability when moving the fiber. By turning the waveplate we noticed that the output power dropped a lot on both polarziation. This behaviour could be due to the fiber connector between the two PM fibers.

- we installed a beam dump on the SL input in order to get rid from possible backreflection when SL unlocks occur.

- we finally lowered the PSTAB gain in order to have the usual UGF of 75 kHz (3rd picture). We had in fact more than 100 kHz.

We relocked everything and we will keep monitoring hte system.

Images attached to this comment
Search Help